is built in the Test Formbox under Banner ID number 1. Nitrogen dioxide was added to a reaction vessel reacts according to
Answer #1

1.

For the given reaction

Kc = N204 [NO2) =  (0.039) (0.4212 = 0.22.

2.

First dissociation constant is highest and third dissocistion constant is lowest.

Ka1 = 7.5×10-3, Ka2 = 6.2×10-8 , Ka3 = 4.8×10-13

3.

When a system is equilibrium

The rate forward and reverse reaction is equal.

Hence option c is correct.

4.

Entropy change when 3 mol oxygen reacted = -139 J.

Or, entropy change when 1 mol oxygen reacted = - (139/3) J

Now, moles of oxygen = (14.6/32) = 0.456

Then entropy change = - (139× 0.456 )/3 = - 21.1 J.

5.

Gibb's energy is negative when a chemical reaction is favourable from left (reactant) to right (product) .